# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

### v2

@@ 4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, we compare our FastVID with state-of-the-art methods on benchmarks with diverse video durations. In DyCoke, video frames are evenly divided into 4-frame segments, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$. For other baselines, we conduct comparisons across different $R$ values.
@@
Table 4: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the retained tokens is from DTM, while the remaining $1-d$ is from ATS.
$d$ in STPrune MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
0.0/ATS 55.3 55.3 62.4 55.0 96.2
0.2 55.0 53.6 62.1 56.1 95.7
0.4 55.9 56.3 62.7 57.3 98.0
0.6 55.6 54.5 62.0 56.2 96.3
0.8 56.0 54.9 62.6 55.6 96.7
1.0/DTM 54.1 50.1 61.5 55.9 93.5
@@
Token Merging MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
@@
Uniform 55.0 54.7 62.0 56.4 96.2
Cluster-based 55.6 55.2 62.4 57.3 97.2
Our DTM 55.9 56.3 62.7 57.3 98.0
@@
Uniform 55.4 54.6 62.9 59.8 91.5
Cluster-based 55.3 55.1 62.9 60.1 91.8
Our DTM 58.5 56.5 64.9 60.7 94.7
@@
[…]
@@ 4.3 Ablation Study
By default, we conduct ablation studies on LLaVA-OneVision at $r=10\%$. Further hyperparameter analysis is provided in the appendix.
@@
Segmentation MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
Fixed-interval 55.1 53.6 61.7 55.4 95.3
Cluster-based 53.2 53.0 61.7 54.8 93.9
Our DySeg 55.9 56.3 62.7 57.3 98.0
@@
Ablation study on STPrune. STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 5 presents an ablation study on $d$, which controls the pruning distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by 1.8% and 4.5% points over ATS and DTM alone, respectively.
@@ 4.4 Qualitative Results
Figure 5 visualizes the proposed DySeg and STPrune. In Figure 5(a), all frames are grouped into a single segment, with the 1st and 3rd frames selected as anchor frames, where DTM is applied. In Figure 5(b), the frames are divided into two segments, with the 1st, 3rd, and 4th frames selected as anchor frames. Blue tokens represent those selected by ATS due to high [CLS] attention. These tokens readily cluster together and do not always correspond to object regions, suggesting that while they reflect salient [CLS] information, they lack broader visual context. In contrast, DTM’s red tokens effectively aggregate visually similar content across frames, significantly reducing spatiotemporal redundancy while preserving visual context. Together, ATS and DTM offer complementary benefits for effective video compression. More qualitative results are provided in the appendix.
@@ 5 Conclusion
  • Section A presents additional experimental settings, including reproduction details of the compared baselines and an estimation of the computational cost.
  • Section B presents additional experimental results, such as ablation studies on key hyperparameters, further qualitative results, and results on Qwen2-VL [37].
  • Section C discusses limitations of our approach and its potential broader impacts.
@@ A.1 Reproduction Details of Compared Baselines
  • VisionZip [^6] https://github.com/dvlab-research/VisionZip, Apache License 2.0 [42] (CVPR 2025). Visionzip performs pruning at the vision encoder’s output, which conflicts with pooling operations in Video LLMs and degrades performance. To address this, we implement VisionZip, which applies pruning after pooling. Following the original settings, each frame retains both dominant and contextual tokens in a $54\!:\!10$ ratio. We define $r$ as the proportion of tokens retained per frame. We set $r\in\{25\%,20\%,15\%,10\%\}$ in Table 1 and $r\in\{25\%,10\%\}$ in Table 2.
  • DyCoke [^7] https://github.com/KD-TAO/DyCoke, Apache License 2.0 [33] (CVPR 2025). DyCoke prunes in both prefilling and decoding stages. For fair comparison, we only evaluate its pre-filling stage. The pruning rate in the TTM module is set $K\in\{0.9,1.0\}$ in both Table 1 and Table 2.
  • PruneVID [^8] https://github.com/Visual-AI/PruneVid, CC BY-NC-SA 4.0 License [13] (ACL 2025). PruneVID includes both input-stage and intra-LLM pruning in the pre-filling phase, along with decoding pruning. For fair comparison, we evaluate two variants: PruneVID (input-stage pruning only) and PruneVID (full prefilling pruning). Following the original settings, we use a threshold $\tau=0.8$, temporal segment ratio $\gamma=0.25$, token selection ratio $\alpha=0.4$, and attention calculations use the 10th layer. The cluster ratio $\beta$ controls compression in the input-stage pruning. We use $\beta=11\%$ in Table 1 and Table 3.
@@ A.2 Computational Cost Estimation
Following prior works [5, 39], we report the theoretical FLOPs of the LLM related to visual/video tokens. Specifically, LLaVA-OneVision [16], LLaVA-Video [47], and Qwen2-VL [37] are all built on Qwen2 [41], which employs grouped-query attention [2] and a three-layer FFN structure. The per-layer FLOPs of the LLM are computed as:
@@ B.2 Ablation study on $p$ and $\beta$ in DTM
B.3 Additional Qualitative Results
@@
Table 7: Comparison of state-of-the-art methods on Qwen2-VL [37]. A/B in # Tokens indicates that A tokens are provided as input to the LLM, and subsequently compressed to B tokens during the LLM forward pass. FastV performs pruning based on LLM attention scores, but its eager-attention implementation materializes the full attention matrix in memory, leading to OOM errors due to the large number of video tokens.
Method # Token TFLOPs VideoMME
Short Medium Long Overall
Vanilla 13447.1 100% 124.04 100% 74.1 60.4 54.3 63.0 100%
FastV [5]${}_{\text{ECCV'24}}$ 13447.1/3361.8 100%/25% 31.34 25.3% Out of Memory
VisionZip [42]${}_{\text{CVPR'25}}$ 3349.3 24.9% 24.11 19.4% 70.9 56.3 48.3 58.5 92.9%
PruneVID [13]${}_{\text{ACL'25}}$ 3460.6 25.7% 24.99 20.1% 66.7 54.0 48.1 56.3 89.4%
FastVID 3349.3 24.9% 24.11 19.4% 72.7 58.3 50.7 60.6 96.2%
B.4 Results on Qwen2-VL
Unlike LLaVA-OneVision [16] and LLaVA-Video [47], Qwen2-VL [37] employs a distinct architecture that samples 768 frames per video and processes them using 3D convolutions. It introduces M-RoPE, which decomposes rotary embeddings into temporal, height, and width components. The model dynamically adjusts the resolution of each frame. In Table 7, our FastVID outperforms other baselines on Qwen2-VL under similar model complexity. In particular, FastVID achieves a 2.4 accuracy improvement on the long subset. As shown in Tables 1, 2, and 7, our FastVID’s consistent superiority across three distinct Video LLM architectures further validates its strong generalizability and practical effectiveness.
@@ Appendix C Additional Discussions
C.1 Limitations
FastVID achieves strong performance on LLaVA-OneVision [16] (32 frames/video), maintaining comparable accuracy with only 8.3% of the FLOPs. However, on LLaVA-Video [47] (64 frames/video) and Qwen2-VL [37] (768 frames/video), although FastVID consistently outperforms existing SoTA baselines, a noticeable accuracy drop occurs compared to the original model. We hypothesize that this is primarily due to the increased number of sampled frames, which leads to fewer tokens per frame and increases the semantic importance of each token. Consequently, aggressive pruning risks discarding critical information. In addition, although our query-agnostic token pruning effectively reduces spatiotemporal redundancy, it fails to emphasize query-relevant information sparsely distributed across extensively sampled frames. To overcome these challenges, future work may explore integrating text-guided keyframe selection to better support long-frame Video LLMs [37, 47].
C.2 Broader Impacts

### v3

@@ 4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, we compare our FastVID with state-of-the-art methods on benchmarks with diverse video durations. We perform evaluations across different retention ratios $R$ to systematically assess model performance. In the case of DyCoke, video frames are evenly divided into 4-frame segments, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$.
@@
Table 6: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the retained tokens is from DTM, while the remaining 1d1-d is from ATS.
$d$
in STPrune
MVBench
LongVideo-
Bench
MLVU VideoMME
Avg.
Acc.
0.0/ATS 55.3 55.3 62.4 55.0 96.2%
0.2 55.0 53.6 62.1 56.1 95.7%
0.4 55.9 56.3 62.7 57.3 98.0%
0.6 55.6 54.5 62.0 56.2 96.3%
0.8 56.0 54.9 62.6 55.6 96.7%
1.0/DTM 54.1 50.1 61.5 55.9 93.5%
@@
Token
Merging
MVBench
LongVideo-
Bench
MLVU VideoMME
Avg.
Acc.
@@
Uniform 55.0 54.7 62.0 56.4 96.2%
Cluster-based 55.6 55.2 62.4 57.3 97.2%
Our DTM 55.9 56.3 62.7 57.3 98.0%
@@
Uniform 55.4 54.6 62.9 59.8 91.5%
Cluster-based 55.3 55.1 62.9 60.1 91.8%
Our DTM 58.5 56.5 64.9 60.7 94.7%
@@
Results on Qwen2-VL. Unlike LLaVA-OneVision [19] and LLaVA-Video [57], Qwen2-VL [45] employs a distinct architecture that samples 768 frames per video and processes them using 3D convolutions. It introduces M-RoPE, which decomposes rotary embeddings into temporal, height, and width components. The model dynamically adjusts the resolution of each frame. In Table 3, our FastVID outperforms other baselines on Qwen2-VL under similar model complexity. In particular, FastVID achieves a 2.4 accuracy improvement on the long subset.
Results on Qwen2.5-VL. To further demonstrate its generalization ability, we report results on Qwen2.5-VL [4] in Table 4. Unlike the LLaVA series and Qwen2-VL, Qwen2.5-VL employs a vision encoder with window attention and a Qwen2.5 LLM [50], representing a substantially different architecture. Compared to the recent SOTA PruneVID, FastVID achieves a 3.5 improvement under a comparable retention ratio. As shown in Tables 1, 2, 3, and 4, our FastVID’s consistent superiority across these distinct Video LLM architectures further validates its strong generalizability and practical effectiveness.
Efficiency Comparison. Table 5 compares the efficiency of our method against the state-of-the-art video compression method PruneVID. Latency and accuracy are evaluated while maintaining similar model complexity. PruneVID relies heavily on time-consuming clustering algorithms during both video segmentation and compression. In contrast, FastVID leverages efficient transition similarity to achieve efficient segmentation. Although density score computation (see Eq. (3-4)) is time-consuming, we restrict this step to anchor frames and parallelize its execution, thereby accelerating the compression process. As a result, FastVID achieves a $\textbf{7.1}\times$ speedup while preserving 98.0% accuracy. Further comparisons with PruneVID are provided in Appendix C.
@@ 4.3 Ablation Study
By default, we conduct ablation studies on LLaVA-OneVision at $r=10\%$. Further hyperparameter analysis is provided in Appendix B.
@@
Segmentation MVBench
LongVideo-
Bench
MLVU VideoMME
Avg.
Acc.
Fixed-interval 55.1 53.6 61.7 55.4 95.3%
Cluster-based 53.2 53.0 61.7 54.8 93.9%
Our DySeg 55.9 56.3 62.7 57.3 98.0%
@@
Ablation study on STPrune. STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 7 presents an ablation study on $d$, which controls the token distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by 1.8% and 4.5% points over ATS and DTM alone, respectively.
@@
Table 9: Ablation study on the SigLIP head in ATS.
Method VideoMME
FastVID 57.3
w/o the SigLIP head 56.3
Table 10: Improved length extrapolation via FastVID.
Method # Frame # Token VideoMME
Vanilla 32 6272 58.6 100%
FastVID (r=25%) 128 6272 60.4 103.1%
FastVID (r=10%) 320 6080 61.4 104.8%
Ablation study on the SigLIP head in ATS. We reintroduce the original, un-finetuned SigLIP head to generate attention scores in ATS. This creates a potential parameter gap between the fine-tuned SigLIP encoder and the original SigLIP head. However, we believe this gap has limited impact for the following reasons.
First, the SigLIP head is pretrained jointly with the encoder on large-scale vision-language data, enabling strong generalization and transferability. Second, since the first-stage training of Video LLMs also aims to align visual and textual representations, the SigLIP head is naturally compatible with this alignment objective. Finally, as shown in Table 10, results show that using the SigLIP head leads to a 1.0 improvement, indicating its effectiveness. This is likely because the SigLIP head is explicitly trained to aggregate patch tokens for semantic alignment, making its attention scores better suited for identifying semantic salient tokens.
In addition, we emphasize that our FastVID generalizes to vision encoders without pretrained heads or [CLS] tokens (e.g., the Qwen-VL series). In such cases, [CLS] attention scores are computed at the Video LLM’s ViT final layer using pseudo [CLS] tokens derived by averaging patch tokens. Our FastVID achieves strong performance on Qwen2-VL (Table 3) and Qwen2.5-VL (Table 4), significantly outperforming previous methods.
Improved length extrapolation. In Table 10, by applying FastVID with different retention ratios, we extend the number of sampled frames from 32 to 320 while maintaining a comparable total number of input video tokens. This consistently improves performance, which demonstrates that FastVID not only compresses video tokens efficiently but also enables effective length extrapolation.
@@ 4.4 Qualitative Results
Figure 5 visualizes the proposed DySeg and STPrune. In Figure 5(a), all frames are grouped into a single segment, with the 1st and 3rd frames selected as anchor frames, where DTM is applied. In Figure 5(b), the frames are divided into two segments, with the 1st, 3rd, and 4th frames selected as anchor frames. Blue tokens represent those selected by ATS due to high [CLS] attention. These tokens readily cluster together and do not always correspond to object regions, suggesting that while they reflect salient [CLS] information, they lack broader visual context. In contrast, DTM’s red tokens effectively aggregate visually similar content across frames, significantly reducing spatiotemporal redundancy while preserving visual context. Together, ATS and DTM offer complementary benefits for effective video compression. More qualitative results are provided in Appendix B.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Natural Science Foundation of China (Nos. 62525103, 62571294), Beijing Natural Science Foundation (No. L252009), CCF-DiDi GAIA Collaborative Research Funds, and the Postdoctoral Science Foundation of China (No. 2024M750565).
@@ Acknowledgments
  • Appendix A presents additional experimental settings, including reproduction details of the compared baselines and an estimation of the computational cost.
  • Appendix B presents additional experimental results, such as ablation studies on key hyperparameters and further quantitative and qualitative results.
  • Appendix C discusses limitations of our approach and its potential broader impacts.
@@ A.1 Reproduction Details of Compared Baselines
  • VisionZip [^6] https://github.com/dvlab-research/VisionZip, Apache License 2.0 [51] (CVPR 2025). Visionzip performs pruning at the vision encoder’s output, which conflicts with pooling operations in Video LLMs and degrades performance. To address this, we implement VisionZip, which applies pruning after pooling. Following the original settings, each frame retains both dominant and contextual tokens in a 54:1054\!:\!10 ratio. We define $r$ as the proportion of tokens retained per frame. We set $r\in\{25\%,20\%,15\%,10\%\}$ in Table 1 and $r\in\{25\%,10\%\}$ in Table 2.
  • DyCoke [^7] https://github.com/KD-TAO/DyCoke, Apache License 2.0 [41] (CVPR 2025). DyCoke prunes during both the prefill and decode stages. For fair comparison, we only evaluate its prefill stage. The pruning rate in the TTM module is set $K\in\{0.9,1.0\}$ in both Table 1 and Table 2.
  • PruneVID [^8] https://github.com/Visual-AI/PruneVid, CC BY-NC-SA 4.0 License [16] (ACL 2025). PruneVID includes both input-stage and intra-LLM pruning in the prefill phase, along with decoding-stage pruning. For fair comparison, we evaluate two variants: PruneVID (input-stage pruning only) and PruneVID (full prefill-stage pruning). Following the original settings, we use a threshold $\tau=0.8$, temporal segment ratio $\gamma=0.25$, token selection ratio $\alpha=0.4$, and attention calculations use the 10th layer. The cluster ratio $\beta$ controls compression in the input-stage pruning. We use $\beta=11\%$ in Table 1 and Table 5.
@@ A.2 Computational Cost Estimation
Following prior works [6, 47], we report the theoretical FLOPs of the LLM related to visual/video tokens. Specifically, LLaVA-OneVision [19], LLaVA-Video [57], and Qwen2-VL [45] are all built on Qwen2 [49], which employs grouped-query attention [2] and a SwiGLU-based three-layer feed-forward network [33]. The per-layer FLOPs of the LLM are computed as:
@@ B.2 Ablation study on $p$ and $\beta$ in DTM
Table 11: Additional quantitative results on VideoMME. We report the overall performance of LLaVA-OneVision.
Method
Retention
Ratio $R$
TFLOPs VideoMME
Vanilla 100% 48.82 58.6
[…]
FastVID 9,7% 4.04 57.3
Table 12: Additional quantitative results on ANet-QA. We report the performance of LLaVA-OneVision.
Method TFLOPs ANet-QA
Accuracy Score
Vanilla 48.82 54.4 3.57
PyramidDrop 14.70 41.2 3.07
FastVID 4.04 53.1 3.52
B.3 Additional Quantitative Results
In Table 12, we conduct additional comparative experiments against previous methods.
Comparison with PyramidDrop [47]. Following its official settings, we prune tokens at the 8th, 16th, and 24th layers with retention ratios of 4.1%, 2.0%, and 1.0%, respectively. Despite aggressive pruning, PyramidDrop still incurs high FLOPs due to full video token processing in the first 8 layers. In contrast, our FastVID compresses tokens from the input stage, reduces computation significantly, and outperforms PyramidDrop under extreme pruning.
Comparison with SparseVLM [56]. Following its official setup, we prune tokens at the 3rd, 7th, and 16th layers with retention ratios of 10.4%, 5.3%, and 2.9%, respectively. Although SparseVLM aggressively prunes later layers, it still incurs high FLOPs due to full token processing in the early layers. In contrast, FastVID compresses tokens from the input stage, leading to significantly lower computational cost. Moreover, FastVID achieves better performance due to its more effective video token pruning strategy.
Comparison with LLaVA-PruMerge [32]. Although both LLaVA-Prumerge and our ATS select tokens based on attention scores, our FastVID is specifically designed to better preserve both temporal and visual context in Video LLMs. A direct quantitative comparison is provided in Table 12. Under the same video token budget of 608 tokens ($\approx$9.7% retention ratio), our method significantly outperforms LLaVA‑PruMerge.
B.4 Results for long-context generation
FastVID prunes redundant video tokens while preserving critical semantic information, which supports both short- and long-context generation tasks. We select ANet-QA [52] for the video captioning task reported in Table 12. Metrics are computed using GPT-3.5-Turbo-0125. FastVID achieves 97.6% accuracy and 98.6% score, demonstrating its effectiveness in long-context generation.
B.5 Additional Qualitative Results
@@
Figures 8(b) and 9(a) showcase failure cases where the query is relevant to only a small subset of frames. In Figure 8(b), only a few frames depict the subject eating a banana, while in Figure 9(a), only a few frames correspond to the entrance scene. In such scenarios, when detailed questions are posed, our query-agnostic pruning strategy becomes less effective. Under high compression rates, most retained tokens are allocated to query-irrelevant frames, making it challenging for the model to provide the necessary visual evidence and produce accurate responses.
@@ Appendix C Additional Discussions
C.1 Comparison with PruneVID
PruneVID [16] and FastVID differ fundamentally in both design and implementation. Below, we detail the key technical distinctions and their practical impact on efficiency and accuracy.
  1. 1.
    Methodological Differences
    1. (a)
      Segmentation Strategy: PruneVID uses the original density-based clustering to partition the video into segments, with a hard upper bound on segment number ($\leq$ sampled frames $\times$ $\gamma$ = 0.25). This constraint may cause semantically diverse frames to be grouped together in complex videos, reducing the effectiveness of subsequent pruning. In contrast, FastVID uses frame transition similarity to adaptively segment the video, ensuring high intra-segment similarity, thereby facilitating more efficient and accurate token pruning.
    2. (b)
      Compression Strategy: PruneVID performs clustering-based merging (Figure 4(a)) on both dynamic and static tokens within each segment, without considering token representativeness or positional structure. In contrast, our proposed density-based token merging DTM selects anchor tokens only from anchor frames and merges the remaining tokens within each segment. The number of anchor tokens per frame is adaptive to segment length. Importantly, our DTM explicitly emphasizes anchor tokens located at density peaks. Positional information of anchor tokens is preserved to maintain the spatiotemporal structure, which is particularly beneficial for the Qwen-VL series that adopts M-RoPE. We further apply anchor-centric aggregation instead of simple average pooling to highlight representative anchor tokens.
  2. 2.
    Practical Impact
    1. (a)
      Efficiency: In Table 5, we conduct a detailed efficiency comparison. The primary bottleneck in efficiency lies in the density score computation. Our segmentation based on frame transition similarity is significantly faster than PruneVID’s clustering. During compression, PruneVID clusters dynamic and static tokens separately per segment. Crucially, the number of tokens varies across segments, which leads to repeated density computations. In contrast, we computes density scores in parallel across all anchor frames, requiring only a single pass. As a result, our pruning speed is $\textbf{6.1}\times$ faster (6.1ms vs. 37.2ms).
    2. (b)
      Accuracy: We compare PruneVID across multiple Video LLMs, including LLaVA-OneVision (Table 1), Qwen2-VL (Table 3), and Qwen2.5-VL (Table 4). Under both the 32 fixed-frame sampling setting (LLaVA-OneVision) and dynamic frame sampling (Qwen-VL series), our method consistently outperforms PruneVID at the similar compression ratio. We attribute this improvement to our more effective dynamic segmentation strategy, which ensures high intra-segment similarity. Furthermore, our token merging strategy highlights representative tokens while retaining thier positional information, making it well-suited for diverse models.
In summary, FastVID introduces a novel framework that integrates dynamic temporal segmentation with density spatiotemporal pruning, effectively preserving both temporal and visual context. It achieves consistent improvements in both efficiency and accuracy over prior methods across diverse Video LLMs.
C.2 Limitations
FastVID achieves strong performance on LLaVA-OneVision [19] (32 frames/video), maintaining comparable accuracy with only 8.3% of the FLOPs. However, on LLaVA-Video [57] (64 frames/video), Qwen2-VL [45] (768 frames/video) and Qwen2.5-VL [4] (768 frames/video), although FastVID consistently outperforms existing SOTA baselines, a noticeable accuracy drop occurs compared to the original model.
This degradation mainly arises from the nature of our query-agnostic pruning strategy. Our aggressive pruning retains only a small number of tokens. When temporal spans are long, few of these retained tokens relate to the query, leading to performance drops. To overcome these challenges, future work may explore integrating query-guided keyframe selection to better support long-frame Video LLMs [45, 57].
However, query-agnostic pruning has a key advantage: the pruned KV cache can be reused across multiple dialogue turns with different queries, making it more suitable for multi-turn conversations. Both query-aware and query-agnostic methods have their own advantages, and the choice can be made based on specific application scenarios.
C.3 Broader Impacts